# A Survey on Large Language Models from General Purpose to Medical Applications: Datasets, Methodologies, and Evaluations

### v1

@@ II-B Data Processing
After acquiring data from the target source, it may not be immediately suitable for training medical LLMs due to several issues: presence of errors or irrelevant information, non-conforming data formats, insufficient or overly brief content, language mismatches between the data source and target, tone inconsistencies with the robotic doctor, and lack of professional annotation. Therefore, this paper summarizes the commonly used data processing methods for medical LLMs and classifies them into four categories according to their functions, namely Data Cleaning, Data Formatting, Data Augmentation, and Translation.
II-B1 Data Cleaning
@@
Data cleaning primarily aims to remove noise and correct errors and inconsistencies, thereby improving overall data quality. We have deeply investigated how the existent medical LLM work cleans the data. This process can be categorized by function into Quality Filtering, Sensitive Information Filtering, and De-duplication.
@@
Quality Filtering enhances data quality by using keywords or special tags to clean the target data. This process can be implemented in various ways, with keyword filtering being a primary method. Keyword filtering helps identify and retain relevant medical domain data, removing non-medical content [97]. By verifying whether a corpus fragment contains pertinent medical keywords, we can effectively determine if it belongs to the desired data. HuatuoGPT-II [45] uses a dictionary-based method to determine if a text fragment belongs to the target domain by evaluating the density of domain-specific words. Similarly, Apollo [54] uses a dictionary-based approach to select medical data from large-scale Chinese, English, and Spanish corpora. MMedLM [49] employs a rule-based filtering pipeline for extracting target data from multilingual corpora. It constructs a list of 1,200 multilingual medical terms and evaluates text passages based on the count and density of these terms, considering a passage relevant only if both metrics exceed a predefined threshold. Additionally, non-essential sections of medical textbooks, such as the cover, table of contents, and end pages, are excluded to maintain a focus on medical content.
@@
Keyword filtering not only facilitates the extraction of target domain data but also enhances data quality. MedAlpaca [29] ensured data quality by selecting only forum replies that received at least five votes. Qibo [55] improved data quality through character-level and paragraph-level rules to verify the plausibility of individual characters and the semantic continuity of the text. These rules were iteratively updated and refined through sampling and manual checks. InMD-X [9] initially identified top-ranking journal titles within each internal medicine subspecialty, based on Journal Citation Reports (JCR), ensuring data quality. Subsequently, it queried these journals via the Pubmed API and extracted papers published since 2010 to capture recent medical knowledge. ChatDoctor [25] improved data quality by filtering out short, meaningless conversations. BianQue [39] reduced data noise by removing private content, incomplete JSON data, website links, voice recordings, and autoresponders from medical consulting platform data using regular expressions. Certain collected data may include medical advertising with misleading language, potentially biasing medical LLMs. HuatuoGPT-II [45] utilizes ChatGPT to identify advertisements within text and trains a quality classification model to filter out such content from the dataset.
@@
The collected data may contain errors, potentially biasing the model’s understanding and leading to inaccuracies in generated text. This could result in misinformation, such as incorrect disease names, which can confuse users. Additionally, grammatical errors in the collected data may be reflected in the model’s output, reducing readability and user experience. To address this, ChatDoctor manually filters erroneous content and utilizes LanguageTool for grammatical corrections, while HuatuoGPT prompts LLMs to filter data for misspellings, slang, and irrelevant information.
@@
The collected data may include special characters like citation marks or URLs, which lack contextual meaning and are unsuitable for training medical LLMs. PMC-LLaMA [30] eliminates irrelevant elements from medical textbooks such as URLs, author lists, redundant information, document content, references, citations, and citation marks for images and tables. MediTron [44] removes URLs, citations, graphics, table separators, and incorrectly formatted characters from clinical guidelines. In processing the literature, author details, bibliographies, acknowledgments, tables, and graphics are excluded, retaining only the main body of each paper. Additionally, it employs “#” to denote major sections and “##” for subsections, replacing citation marks with the title of the reference.
@@
Although manual review is time-consuming and labor-intensive, it remains the most reliable method for ensuring data quality. OncoGPT [53] submits automatically cleaned data for review by domain experts and clinicians to correct inaccuracies and ensure the authority of the answers.
@@
Sensitive Information Filtering. Private information in medical data usually refers to patients’ personal identity and health information, which can seriously harm patients’ rights if disclosed. Moreover, including private attributes like race, gender, and age may introduce bias in certain demographic groups. Common techniques for mitigating privacy concerns include data desensitization, substituting personal identifiers for sensitive information within the text, and data deletion, which involves directly removing sensitive details from the text. However, the latter approach may compromise the integrity and usability of the data. ChiMed-GPT and OncoGPT remove private information, including personal details, hospital data, and website information, from dialog data. This process mitigates the impact of data redundancy and non-critical information on model performance.
@@
De-duplication removes duplicates, reduces data redundancy, and enhances the quality and diversity of training data. Additionally, it prevents data leakage between the training and test sets, ensuring a fair evaluation process. HuatuoGPT-II [45] employs a sentence embedding model to transform the corpus into embedding vectors and subsequently eliminates semantically similar texts through an intensive search method. MediTron [44] de-duplicates literature samples by matching paper titles. Apollo [54] performs de-duplication on exam data by removing samples with at least 64 consecutive identical characters to prevent data leakage and ensure accurate assessment results.
II-B2 Data Formatting
@@
The continued pretraining phase requires unstructured data, the instruction fine-tuning phase requires QA pairs (both single-round and multi-round), and the human alignment phase involves human preference data, where preferences are labeled on top of the QA pairs. The collected data may not be directly usable for the training stages, necessitating transformation into the appropriate format. We summarize the methods used by medical LLMs work to convert data from various structures into three types: Unstructured Data, Instruction Data, and Human Preference Data.
@@
Unstructured Data sare sourced from medical encyclopedias, textbooks, academic literature, and guidelines. These data are used in the continued pretraining phase of medical LLMs to enrich domain knowledge and enhance representation capabilities. After a simple truncation process, these unstructured data can effectively support the model’s continued pretraining. However, structured data needs to be transformed into unstructured data for continued pretraining. For knowledge graph data, entities, relationships, and attributes are extracted. Subsequently, these information can be transformed into unstructured natural language text using human-designed templates or natural language generation tools like LLM. AntGLM-Med [46] samples subgraphs at different scales of the knowledge graph and rewrites them as natural language text. Questio-answer pairs and NLP task data are commonly stored in JSON format. This format facilitates the extraction of data and labels, enabling the formation of unstructured natural language text using key-value pairs as delimiters. MedicalGPT [35] aggregates questions and answers from the medical encyclopedia QA dataset to create plain text data for continued pretraining, thereby integrating medical knowledge.
@@
Instruction Data consists of question-and-answer formats, facilitating medical LLMs in comprehending the context and concepts of medical queries. It also aids in familiarizing the models with common expressions and language conventions in the medical domain, resulting in more natural and clinically relevant responses. Both synthetic data and real doctor-patient conversation data adhere to the question-answer pair format. However, for sources like medical literature, encyclopedias, textbooks, and knowledge graphs, transformation into question-answer pairs requires the application of natural language generation techniques. MedAlpaca [29] uses ChatGPT to transform paragraph headings from medical textbooks into questions and paragraph content into corresponding answers, thereby generating question-answer pair dataset. Apollo [54] partitions the corpus data into segments based on basic semantic units, such as chapters in books, paragraphs in website content, and abstracts of papers. Subsequently, it prompts ChatGPT to sequentially generate questions and answers based on these text segments. InMD-X [9] prompts ChatGPT to extract fundamental medical knowledge from abstracts of medical literature. It subsequently prompts ChatGPT to generate corresponding questions based on these basic medical knowledge. Me-LLaMA [50] formulates specific prompts for various datasets, including medical literature, clinical notes, clinical guides, Wikipedia, and Knowledge Graphs. It then prompts the LLM to enhance the original questions, thereby creating instruction fine-tuning datasets. Similarly, HuatuoGPT-II [45] prompts ChatGPT to convert the unstructured medical corpus into a question-answer pair format. Additionally, it introduces statistical analysis and semantic recognition techniques to minimize discrepancies between the transformed data and the original source. Clinical Camel [33] presents a dialogue-based knowledge encoding approach, which converts dense medical literature into dialogues and incorporates soft alignment. Initially, it employs a teacher model to convert a segment of literature text into a multi-round dialogue, with alignment conditions included in the prompts. Subsequently, a student model is tasked with learning the mapping between the literature text and the multi-round dialogue, resulting in an enhanced student model that improves conversational capabilities. BiMediX [52] utilizes samples from a multiple-choice question-and-answer dataset and specific prompt words as inputs to ChatGPT, generating a multi-round question-and-answer dataset. DISC-MdeLLM [37] employs LLM to distill questions and correct answers from the multiple-choice question-and-answer dataset. It then integrates these with explanations to produce question-and-answer-formatted data.
@@
The above work employs LLMs for data formatting, while other work leverage automated methods for the same purpose. ClinicalGPT utilizes annotations from electronic health records as input for instruction fine-tuning data, with diagnoses as output, forming a dataset of question-answer pairs. GPT-doctor utilizes package inserts to generate QA pairs covering aspects such as usage, dosage, indications, contraindications, precautions, pharmacological effects, chemical composition, and more. The questions are constructed based on subheadings in the package inserts, like “What diseases does [Medicine Name] treat?”, and then the answer is the corresponding content of the subheading.
@@
The medical knowledge graph embodies a wealth of structured medical knowledge. Efforts in medical LLMs involve the conversion of knowledge graph data into QA pairs format. BenTsao [27] first extracts medical knowledge described in natural language from the knowledge graph, and then generates the corresponding instructions and answers based on this knowledge using ChatGPT. DISC-MedLLM [37] initially prompts ChatGPT to convert disease-related knowledge into the $<$instruction, knowledge$>$ format, and then prompts ChatGPT to convert these data into a single-round medical dialogue scenario $<$user inquiry, physician response$>$. This process aims to augment the diversity and linguistic depth of training datasets. ClinicalGPT [36] manually designed the corresponding templates to convert each triple from the knowledge graph into data in QA pair format. Likewise, PMC-LLaMA [30] formulated QA templates to either describe an entity or predict the relationship between two entities.
@@
Human Preference Data is intended to be used to train the reward model and to guide the model in aligning human preferences such as being more doctor-like, patient-friendly, truthful, harmless, and professional. This type parallels the QA format, where responses are categorized into chosen and rejected, based on their alignment with the predefined human preferences. The construction of the human preference dataset is more time-consuming compared to the instruction fine-tuning dataset, Therefore, many studies leverage ChatGPT to facilitate the construction of the human preference dataset. HuatuoGPT [8] utilizes real dialogues to stimulate the generative model into producing diverse responses. These responses undergo evaluation via ChatGPT, considering coherence, consistency with human preferences, and comparison with responses from real doctors. The resultant human preference dataset is structured based on the ratings provided by ChatGPT. ChiMed-GPT [43] and MedicalGPT [35] sample questions from the instruction fine-tuning dataset, then the chosen response is provided by the physician and the rejected response is provided by the medical LLM BenTsao [27]. To pursue high-quality human preference data, some of the work uses manual annotation. Zhongjing [38] employs medical graduate students or clinicians to annotate the human preference dataset with nine dimensions (Accuracy, Safety, Ethics, Comprehension, Clarity, Initiative, Coherence, Consistency, and Warm Tone), yielding a dataset of superior quality. Similarly, ClinicalGPT [36] employs manual annotation to rank responses to medical inquiries.
II-B3 Data Augmentation
@@
Short, colloquial responses from doctors and multiple-choice QA datasets lacking explanations in the collected conversation data may compromise the accuracy and professionalism of medical LLMs. Consequently, some studies use LLMs to augment the training dataset for addressing these issues. BianQue [39] prompts ChatGPT to refine doctors’ responses over multiple rounds of conversations, enhancing their empathy, professionalism, and coherence. DISC-MedLLM [37] prompts ChatGPT to expand on the original responses, adding detail and logic without altering their meaning. It also eliminates colloquial expressions and revises responses that do not align with the style of the robot doctor, such as advising patients to make appointments. Zhongjing [38] employs a self-instruct [96] method to standardize, professionalize, and enhance the friendliness of doctors’ responses while preserving the diversity and integrity of the original data. PMC-LLaMA [30] prompts GPT-4 to augment medical conversation data by generating synonymous sentences, thereby improving the model’s robustness to different instructions. Additionally, it uses ChatGPT to add rationale explanations to multiple-choice QA data. Clinical Camel [33] prompts the GPT-4 to retrieve textual resources related to the question, and then supplements the multiple-choice QA data with answer explanations according to those resources. Some studies have used LLMs to augment human preference datasets. ChiMed-GPT [43] enhanced the human preference dataset by incorporating responses from GPT-4 and GPT-3.5. The response format changed from $<$chosen response, rejected response$>$ to $<$chosen response, GPT-4’s response, GPT-3.5’s response, rejected response$>$, with the degree of alignment to human preferences decreasing in that order.
II-B4 Translation

### v2

@@ 2.2 Data Processing
After acquiring data from the target source, it may not be immediately suitable for training medical LLMs due to several issues: the presence of errors or irrelevant information, non-standardized data formats, insufficient or overly brief content, language mismatches between the data source and target, tone inconsistencies with the robotic doctor, and lack of professional annotation. Therefore, this survey summarizes the commonly used data processing methods for medical LLMs and classifies them into four categories according to their functions: Data Cleaning, Data Formatting, Data Augmentation, and Translation. Additionally, we summarized the related prompt templates of data processing in the GitHub repository.
2.2.1 Data Cleaning
@@
Data cleaning primarily aims to remove noise and correct errors and inconsistencies, improving overall data quality [27]. We have deeply investigated how existing medical LLM works to clean these data. This process can be categorized by function into Quality Filtering, Sensitive Information Filtering, and De-duplication.
@@
Quality Filtering enhances data quality by using keywords or special tags to clean the target data. This process can be implemented in various ways, with keyword filtering being a primary method. Keyword filtering helps identify and retain relevant medical domain data, removing non-medical content [168]. By verifying the presence of pertinent medical keywords in a corpus fragment, we can effectively determine its relevance to the desired data. HuatuoGPT-II [22] uses a dictionary-based method to determine if a text fragment belongs to the target domain by evaluating the density of domain-specific words. Similarly, Apollo [156] uses a dictionary-based approach to select medical data from large-scale Chinese, English, and Spanish corpora. MMedLM [121] employs a rule-based filtering pipeline for extracting target data from multilingual corpora. It constructs a list of 1,200 multilingual medical terms and evaluates text passages based on the count and density of these terms, considering a passage relevant only if both metrics exceed a predefined threshold. Additionally, non-essential sections of medical textbooks, such as the cover, table of contents, and end pages, are excluded to maintain a focus on medical content.
@@
Keyword filtering not only facilitates the extraction of target domain data but also enhances data quality. MedAlpaca [54] ensured data quality by selecting only forum replies that received at least five votes. Qibo [189] improved data quality through character-level and paragraph-level rules to verify the plausibility of individual characters and the semantic continuity of the text. These rules were iteratively updated and refined through sampling and manual checks. InMD-X [53] initially identified top-ranking journal titles within each internal medicine specialty, based on Journal Citation Reports (JCR), ensuring data quality. Subsequently, it queried these journals via the Pubmed API and extracted papers published since 2010 to capture recent medical knowledge. ChatDoctor [91] improved data quality by filtering out short, meaningless conversations. BianQue [23] reduced data noise by removing private content, incomplete JSON data, website links, voice recordings, and auto-responds from medical consulting platform data using regular expressions. Additionally, collected data may include medical advertising with misleading language, potentially biasing medical LLMs. HuatuoGPT-II [22] utilizes ChatGPT to identify advertisements within text and trains a quality classification model to filter out such content from the dataset.
@@
The collected data may contain errors, potentially biasing the model’s understanding and leading to inaccuracies in the generated text. This could result in misinformation, such as incorrect disease names, which can confuse users. Additionally, grammatical errors in the collected data may be reflected in the model’s output, reducing readability and user experience. To address this, ChatDoctor [91] manually filters erroneous content and utilizes LanguageTool for grammatical corrections, while HuatuoGPT [188] prompts LLMs to filter data for misspellings, slang, and irrelevant information.
@@
The collected data may include special characters like citation marks or URLs, which lack contextual meaning and are unsuitable for training medical LLMs. PMC-LLaMA [167] eliminates irrelevant elements from medical textbooks such as URLs, author lists, redundant information, document content, references, citations, and citation marks for images and tables. MediTron [24] removes URLs, citations, graphics, table separators, and incorrectly formatted characters from clinical guidelines. In processing the literature, author details, bibliographies, acknowledgments, tables, and graphics are excluded, retaining only the main body of each paper. Additionally, it employs “#" to denote major sections and “##" for subsections, replacing citation marks with the title of the reference.
@@
Although the manual review is time-consuming and labor-intensive, it remains the most reliable method for ensuring data quality. OncoGPT [69] submits automatically cleaned data for review by domain experts and clinicians to correct inaccuracies and ensure the authority of the answers.
@@
Sensitive Information Filtering. Private information in medical data usually refers to patients’ personal identity and health information, which can seriously harm patients’ rights if disclosed [47, 101]. Moreover, including private attributes like race, gender, and age may introduce bias in certain demographic groups. Common techniques for mitigating privacy concerns include data desensitization, substituting personal identifiers for sensitive information within the text, and data deletion, which involves directly removing sensitive details from the text. However, the latter approach may compromise the integrity and usability of the data. ChiMed-GPT [142] and OncoGPT [69] remove private information, including personal details, hospital data, and website information, from dialogue data. This process mitigates the impact of data redundancy and non-critical information on model performance.
@@
De-duplication removes duplicates, reduces data redundancy, and enhances the quality and diversity of training data. Additionally, it prevents data leakage between the training and test sets, ensuring a fair evaluation process [14, 44]. HuatuoGPT-II [22] employs a sentence embedding model to transform the corpus into embedding vectors and subsequently eliminates semantically similar texts through an intensive search method. MediTron [24] de-duplicates literature samples by matching paper titles. Apollo [156] performs de-duplication on exam data by removing samples with at least 64 consecutive identical characters to prevent data leakage and ensure accurate assessment results.
2.2.2 Data Formatting
@@
The continued pretraining stage requires unstructured data, the instruction fine-tuning stage requires QA pairs (single-turn and multi-turn), and the human alignment stage involves human preference data. The collected data may not be directly suitable for the training stages, necessitating transformation into the appropriate format. We summarize the methods used by medical LLMs to convert data from various structures into three types: Unstructured Data, Instruction Data, and Human Preference Data.
@@
Unstructured Data are sourced from medical encyclopedias, textbooks, academic literature, and guidelines. These data are used in the continued pretraining stage of medical LLMs to enrich domain knowledge and enhance representation capabilities. After a simple truncation process, these unstructured data can effectively support the model’s continued pretraining. However, structured data needs to be transformed into unstructured data for continued pretraining. For knowledge graph data, entities, relationships, and attributes are extracted. Subsequently, these information can be transformed into unstructured natural language text using human-designed templates or natural language generation tools like LLM. AntGLM-Med [89] samples subgraphs at different scales of the knowledge graph and rewrites them as natural language text. QA pairs and NLP task data are commonly stored in JSON format. This format facilitates the extraction of data and labels, enabling the formation of unstructured natural language text using key-value pairs as delimiters. MedicalGPT [174] aggregates questions and answers from the medical encyclopedia QA dataset to create plain text data for continued pretraining, thereby integrating medical knowledge.
@@
Instruction Data consists of question-and-answer(QA) formats, facilitating medical LLMs in comprehending the context and concepts of medical queries. It also aids in familiarizing the models with common expressions and language conventions in the medical domain, resulting in more natural and clinically relevant responses [42, 114]. Both synthetic data and real doctor-patient conversation data adhere to the question-answer pair format. However, for sources like medical literature, encyclopedias, textbooks, and knowledge graphs, transformation into question-answer pairs requires the application of natural language generation techniques. MedAlpaca [54] uses ChatGPT to transform paragraph headings from medical textbooks into questions and paragraph content into corresponding answers, thereby generating QA pair dataset. Apollo [156] partitions the corpus data into segments based on basic semantic units, such as chapters in books, paragraphs in website content, and abstracts of papers. Subsequently, it prompts ChatGPT to sequentially generate questions and answers based on these text segments. InMD-X [53] prompts ChatGPT to extract fundamental medical knowledge from abstracts of medical literature. It subsequently prompts ChatGPT to generate corresponding questions based on these basic medical knowledge. Me-LLaMA [171] formulates specific prompts for various datasets, including medical literature, clinical notes, clinical guides, Wikipedia, and Knowledge Graphs. It then prompts the LLM to enhance the original questions, thereby creating instruction fine-tuning datasets. Similarly, HuatuoGPT-II [22] prompts ChatGPT to convert the unstructured medical corpus into a question-answer pair format. Additionally, it introduces statistical analysis and semantic recognition techniques to minimize discrepancies between the transformed data and the original source. Clinical Camel [144] presents a dialogue-based knowledge encoding approach, which converts dense medical literature into dialogues and incorporates soft alignment. Initially, it employs a teacher model to convert a segment of literature text into a multi-turn dialogue, with alignment conditions included in the prompts. Subsequently, a student model is tasked with learning the mapping between the literature text and the multi-turn dialogue, resulting in an enhanced student model that improves conversational capabilities. BiMediX [118] utilizes samples from a multiple-choice QA dataset and specific prompt words as inputs to ChatGPT, generating a multi-turn QA dataset. DISC-MdeLLM [11] employs LLMs to distill questions and correct answers from the multiple-choice QA dataset. It then integrates these with explanations to produce single-turn QA data.
@@
The above work employs LLMs to perform data formatting, while other works leverage automated methods for the same purpose. ClinicalGPT [150] utilizes annotations from electronic health records as input for instruction fine-tuning data, with diagnoses as output, forming a dataset of question-answer pairs. GPT-doctor [155] utilizes package inserts to generate QA pairs covering aspects such as usage, dosage, indications, contraindications, precautions, pharmacological effects, chemical composition, and more. The questions are constructed based on subheadings in the package inserts, like “What diseases does [Medicine Name] treat?", and then the answer is the corresponding content of the subheading.
@@
The medical knowledge graph embodies a wealth of structured medical knowledge. Efforts in medical LLMs involve the conversion of knowledge graph data into QA pairs format. BenTsao [151] first extracts medical knowledge described in natural language from the knowledge graph, and then generates the corresponding instructions and answers based on this knowledge using ChatGPT. DISC-MedLLM [11] initially prompts ChatGPT to convert disease-related knowledge into the $<$instruction, knowledge$>$ format, and then prompts ChatGPT to convert these data into a single-turn medical dialogue scenario $<$user inquiry, physician response$>$. This process aims to augment the diversity and linguistic depth of training datasets. ClinicalGPT [150] manually designed the corresponding templates to convert each triple from the knowledge graph into data in QA pair format. Likewise, PMC-LLaMA [167] formulated QA templates to either describe an entity or predict the relationship between two entities.
@@
Human Preference Data is intended to be used to train the reward model and to guide the LLM in aligning human preferences such as being more doctor-like, patient-friendly, truthful, harmless, and professional [133, 188]. This type parallels the QA format, where responses are categorized into chosen and rejected, based on their alignment with the predefined human preferences. The construction of the human preference dataset is more time-consuming compared to the instruction fine-tuning dataset, Therefore, many studies leverage ChatGPT to facilitate the construction of the human preference dataset. HuatuoGPT [188] utilizes real dialogues to stimulate the generative model into producing diverse responses. These responses undergo evaluation via ChatGPT, considering coherence, consistency with human preferences, and comparison with responses from real doctors. The resultant human preference dataset is structured based on the ratings provided by ChatGPT. ChiMed-GPT [142] and MedicalGPT [174] sample questions from the instruction fine-tuning dataset, then the chosen response is provided by the physician, and the rejected response is provided by the medical LLM BenTsao [151]. To pursue high-quality human preference data, some of the work uses manual annotation. Zhongjing [179] employs medical graduate students or clinicians to annotate the human preference dataset with nine dimensions (Accuracy, Safety, Ethics, Comprehension, Clarity, Initiative, Coherence, Consistency, and Warm Tone), yielding a dataset of superior quality. Similarly, ClinicalGPT [150] employs manual annotation to rank responses to medical inquiries.
2.2.3 Data Augmentation
@@
Shorter, colloquial doctor-patient conversation data and multiple-choice QA datasets lacking detailed explanations may compromise the accuracy and professionalism of responses generated by medical LLMs. Consequently, some studies use LLMs to augment the training dataset for addressing these issues. BianQue [23] prompts ChatGPT to refine doctors’ responses over multiple rounds of conversations, enhancing their empathy, professionalism, and coherence. DISC-MedLLM [11] prompts ChatGPT to expand on the original responses, adding detail and logic without altering their meaning. It also eliminates colloquial expressions and revises responses that do not align with the style of the robot doctor, such as advising patients to make appointments. Zhongjing [179] employs a self-instruct [160] method to standardize, professionalize, and enhance the friendliness of doctors’ responses while preserving the diversity and integrity of the original data. PMC-LLaMA [167] prompts GPT-4 to augment medical conversation data by generating synonymous sentences, thereby improving the model’s robustness to different instructions. Additionally, it uses ChatGPT to add rationale explanations to multiple-choice QA data. Clinical Camel [144] prompts the GPT-4 to retrieve textual resources related to the question, and then supplements the multiple-choice QA data with answer explanations according to those resources. Some studies have used LLMs to augment human preference data. ChiMed-GPT [142] enhanced the human preference dataset by incorporating responses from GPT-4 and GPT-3.5. The response format changed from $<$chosen response, rejected response$>$ to $<$chosen response, GPT-4’s response, GPT-3.5’s response, rejected response$>$, with the degree of alignment to human preferences decreasing in that order.
2.2.4 Translation